OpenStudy (anonymous):

How would you work this problem? Working alone, Ryan can dig a 10ft by 10ft hole in five hours. Castel can dig the same hole in six hours. How long would it take them if they worked together?

OpenStudy (stamp):

\[Ryan=\frac{5\ hours}{1\ hole}\]\[Castel=\frac{6\ hours}{1\ hole}\]Does this help you find an approach? What answer do you obtain for t hours?

OpenStudy (anonymous):

All I see is dividing their time by the number of holes, (1 hole) and that would still be the number I started with. So I still don't understand.

OpenStudy (stamp):

\[Ryan+Castel=\frac{5\ hours}{1\ hole}+\frac{6\ hours}{1\ hole}\]\[=\frac{11\ hours}{2\ holes}=6.5\ hours\]I will leave it to you to check the answer

OpenStudy (anonymous):

Oh, okay. Now I get it. Thank you :)